I need help with this word problem please.
On three consecutive passes, a football team gains 9 yards, loses 26yards and gains 28yards.
What number represents the total net yardage?
The total net yardage is ? yards.

Answer by Earlsdon(6103)

You need to add the following integers: (Gain = +) (Loss = -)
9+(-26)+28 = 9+2 = 11
11 yards is the total net gain.